# AFaCTA: Assisting the Annotation of Factual Claim Detection with Reliable LLM Annotators

### v1

@@ Appendix I Self-Consistency CoT
Error 2: When our federal government overreached, we found a way to fight back. Error reason: the model argues it lacks details of “overreach”. Error 3: While our work is far from over, we have made significant progress thanks to the Rebuild Alabama Act. Error reason: the model argues it lacks important details of “significant progress”. Error 4: Folks, no doubt, the last couple of years have been especially trying for our medical professionals. Error reason: the model argues the “especially trying” lacks detail. Error 5: I am proud that my Administration, with the support of the Legislature, is doing more to make significant improvements in mental health care than any since Governor Lurleen Wallace in the 1960s. Error reason: the model argues the “significant improvements” lack detail. Error 6: At times, her schoolwork and distance from her home state made her wonder if she should give up her Miss Alaska title. Error reason: the model argues the “significant improvements” lack detail. Error 7: It’s the result of great investment decisions, policies, vision, and direction. Error reason: not enough context about “it”. Error 8: Together with these partners, we’ll build a stronger, more durable health care system in Alaska that can respond to most any situation. Error reason: the model argues it lacks details about the plan. Error 9: At the same time, our ability to increase production is under attack from Washington, DC, and federal courts that side with extremist environmental groups. Error reason: the model argues it lacks details about the “attack”. Error 10: No state has been targeted more by the current administration than our Great State of Alaska. Error reason: the model argues it lacks details about specific actions. Error 11: No state has been targeted more by the current administration than our Great State of Alaska. Error reason: the model argues it lacks details about specific actions. Error 12: At every turn and since day one of the Biden Administration, this hostility has been perfectly clear. Error reason: the model argues it lacks details about the “hostility”. Error 13: Because no president should have to beg for more oil from the Middle East or Russia’s Arctic when we can produce it right here better and safer than anywhere else on the planet! This is common sense! Error reason: the model argues it lacks details or evidence about “the US produces better oil”. Error 13: Many of them have been with us for so long that they’ve almost been normalized in Alaska, as almost unsolvable. Error reason: lacking context and details about “long issues”. Error 14: I will always stand between Alaskans and a federal government that violates our God-given rights and exceeds its constitutional authority. Error reason: the model argues it lacks details about specific actions of the federal government and the speaker’s future action. Error 15: I envision an Alaska where our cost of energy is no longer the second-highest in the nation, but one of the lowest. That’s my vision. I hope it is yours as well. Error reason: the model argues it lacks details about the definition of “second-highest” and “lowest”. Error 16: I’ve seen it in the men and women on the frontlines of this pandemic who have helped us achieve one of the shortest shutdowns and one of the lowest death rates in the country. Error reason: the model argues it lacks details about “shortest” and “lowest”. Error 17: And because we want to lead by example, we are saving Coloradans money by making your State Government more efficient and effective. Error reason: the model argues it lacks details about “efficient” and “effective”. Error 18: Just as an earthquake is followed by aftershocks, we know that the overarching crisis of the pandemic has led to many other crises, perhaps lesser seen, but no less important to address. Error reason: the model argues it lacks details about “the crises”. Error 19: We owe it to the people of Colorado to improve safety and make Colorado truly one of the ten safest states in the nation over the next five years. Error reason: the model argues it lacks details about the speaker’s plan. Error 20: No other place offers opportunity to so many from such diverse backgrounds. Error reason: the model argues it lacks specific details. Error 21: It’s that, as our businesses grow, we don’t leave our workers behind. Error reason: the model argues it lacks specific details about business growth. Error 22: By creating choices - real choices - for parents, and unprecedented support for their kids. Error reason: the model argues it lacks specific details about the choices and supports. |
@@
False Negative: understand facts as opinions or fail to identify facts entangled with opinions |
Error 23: They plan their lives around hunting season, or fishing season; construction season, or tourism season.But not election season. Error reason: the model misunderstands it as the speaker’s opinion. But people’s lifestyles and priorities can be verified with related surveys or studies. Error 24: A future where a dynamic, multi-modal transportation system meets the needs of our growing population. Error reason: the model fails to identify “our growing population” as a fact. Error 25: When I was elected Governor, I knew that I would be remembered not for who I was, where I came from, or even what I said at events like this, but for what I did to make a meaningful, measurable, positive impact on the lives of Coloradans. Error reason: the model fails to identify that “the speaker is elected as the governor” is a fact. Error 26: But over time, we’ve learned we can’t solve big problems like climate change situationally, with short-term thinking. Error reason: the model fails to identify the causality claim about short-term thinking and big problems. Error 27: But at a time, when we’ve been heating and burning up, one thing we cannot do is repeat the mistakes of the past by embracing polluters. Error reason: the model fails to recognize the fact of embracing polluters in the past. |
False Negative: hallucinate about personal experience and citation |
Error 28: At times, her schoolwork and distance from her home state made her wonder if she should give up her Miss Alaska title. Error reason: the model argues the personal experience is subjective. Error 29: “A lot of people,” she said, “don’t recognize that their low points are what are going to propel them to their future. Error reason: subjective personal experience. Error 30: I agree with former Governor Jay Hammond that the government should never take more from the Permanent Fund than is distributed to the people of Alaska. Error reason: fail to detect the citation. Error 31: She is in a healthy marriage and is reconnecting with her children. Error reason: consider personal experience as unverifiable. Error 32: “Dad,” Catherine said, “Alaska has so much to offer.” Error reason: fail to detect the citation. Error 33: Still, she found the strength to take down the shooter, ending his violent killing spree and saving many precious lives. Error reason: consider personal experience as unverifiable. |
False Negative: hallucinate about rhetoric |
Error 34: They’re wondering how we’ve come to a place where the PFD is nothing more than what’s left over after government takes the lion’s share. Error reason: fail to understand the metaphor. |
@@
False Positive: over-sensitive to granular, unspecific, or not-explicitly-presented facts |
Error 1: Requesting to work from home because of the #coronavirus is what’s called a “reasonable accommodation.” You have disabled people to thank for that. Remember this moment in history the next time you think Accessibility laws are too “burdensome” to be abided. Error reason: the model recognizes the concept of “reasonable accommondation” and the existence of “accessibility laws” as facts, which are not explicitly presented by the post. |
False Negative: misunderstand verifiable fact as subjective interpretation |
Error 2: “Last week Trump told aides he’s afraid journalists will try to purposefully contract #coronavirus to give it to him on Air Force One.” https://t.co/sS1MZR6D7w Error reason: GPT-4 understands it as the tweet author’s subjective interpretation of Trump’s words. However, we think that it can be verified by checking whether Trump said the words or not. Error 3: Due to #coronavirus, media advises the economy must tank, the people must panic, Trump must be blamed, Biden must be secreted away from the public, and Bernie must cease rallies. I wonder why people do not trust the media’s motives on this? Error reason: GPT-4 understands it as the tweet author’s subjective interpretation of the media’s advice. However, we think it can be verified by checking if there are media suggesting such information. |
@@
False Negative: fail to identify facts entangled with opinions |
@@
Error 1: Who would you prefer to lead our nation’s response to the growing #coronavirus threat? Error reason: fail to identify “the growing coronavirus threat”. Error 2: It was a really really really really really really really really really really really really really really really really really really really really bad idea to elect Donald Trump President of the United States. #TrumpVirus #TrumpCrash #TrumpRecession #COVID19 #coronavirus Error reason: fail to identify “elected Donald Trump Predisent of the US”. Error 3: If people who are infected by corona virus in SA were black, their names, homes street: pictures will be all over Social Media. White privileges goes a long way. Wait for a case of a black person, they will mention even his location they won’t say WC, they’ll say Gugulethu ext 5 Error reason: fail to identify the correlation between the infected persons’ race and their suffers. Error 4: @realDonaldTrump On a morning when Americans are terrified, the markets are gonna historically crash and we need LEADERSHIP…all you’ve done is hate-tweet BULLSHIT about Sanders, Warren, Biden, Democrats, Schumer. the media and now Obama. Your incompetence is staggering…. #Trump #coronavirus Error reason: although have subjective interpretations, the facts that “markets are gonna historically crash” and Trump commented something about others is verifiable. Error 5: Dear BBC, I want to fight for you. I know you’re more than news (which has been questionable) you’re also great drama, documentaries, kids tv etc But don’t make me question that by inviting Farage on to talk about Corona FFS!! Show you’ll fight for your integrity yourselves! Error reason: fail to identify the verifiable fact that BBC invited Farage is verifiable. Error 6: Thread 1: One wonders about the racial politics of this corona outbreak. What would have happened had it been blacks who came into the country with the virus? Would they hav been allowed to “self quarantine”? If the virus was from the continent; wouldn’t travels be banned by now? Error reason: fail to identify “this corona outbreak” and the practice of “self-quarantine”. Error 7: The total Iranian #COVID19 case-count is in the hundreds of thousands, perhaps millions, according to my estimates (detailed at the link). This raises an important question: if there are two million cases, where are all the bodies? https://t.co/nHYbQlXlVC Error reason: GPT-3.5 understands it as the tweet author’s subjective interpretation about the numbers. However, it can be verified by checking details in the link and reliable data source. Error 8: Due to #coronavirus, media advises the economy must tank, the people must panic, Trump must be blamed, Biden must be secreted away from the public, and Bernie must cease rallies. I wonder why people don’t trust the media’s motives on this? Error reason: GPT-3.5 understands it as the tweet author’s subjective interpretation about the media’s advice. However, we think it can be verified by checking if there are medias suggesting such information. |
False Negative: fail to comprehend claims about attached links |
@@
Error 9: Public Safety Announcement Fighting #CoronaVirus. We have to do this together. Wishing good health to all of you! Love, Vijay. https://t.co/fbafmmtq8S Error reason: this tweet claims that the link contains a public safetly announcement fighting COVID, which is verifiable. Error 10: This thread needs to fly. It shows how the legacy media is USING covid-19 as a political weapon and even how the SAME reporters are contradicting themselves. This. Is. SICK. https://t.co/Werq544xii Error reason: this tweet claims the content of the link, which is verifiable. |
False Negative: fail to identify personal experience or citation |
@@
Error 11: Beware the spread of coronavirus and Fox News pandemic propaganda. Trish Regan’s melodramatic rant decrying Dems and MSM for allegedly exploiting #COVID19 as “another attempt to impeach the President.” Yank this dangerous shrew off the air. #Trumpdemic https://t.co/6B60RLMIS0 Error reason: fail to identify the personal experience of Trish Regan. Error 12: I keep bumping into this problem. I want to be able to stand up and unequivocally defend the BBC. But it has repeatedly shut out radical voices and crucial issues while providing a massive platform for the alt-right to spout ill-informed nonsense. It is hard to love. https://t.co/A1pQMsqDxV Error reason: The interpretation about BBC’s behavior is subjective. But the tweet author’s previous stance might be verifiable by checking their previous statements. Error 13: @keywilliamss One African man actually got Corona and was cured in like a week. Health care officials are “baffled as to why Africa is virtually unscathed.” Which is… kinda racist that they expected it to be but lemme hush URL: https://t.co/yvP0DUXDiX Error reason: fail to identify the African man’s experience and the quotation of health care officials’ speech. |
@@
We conduct a thorough analysis on GPT-4 and GPT-3.5 AFaCTA. Errors on PoliClaim can be found in Table 8, Table 9, and Table 10. Errors on CheckThat!-2021-dev can be found in Table 11 and Table 12.

### v3

@@ Appendix I Self-Consistency CoT
| False Negative: not enough specific detail or context and thus not verifiable | Error 2: When our federal government overreached, we found a way to fight back. Error reason: the model argues it lacks details of “overreach”. Error 3: While our work is far from over, we have made significant progress thanks to the Rebuild Alabama Act. Error reason: the model argues it lacks important details of “significant progress”. Error 4: Folks, no doubt, the last couple of years have been especially trying for our medical professionals. Error reason: the model argues the “especially trying” lacks detail. Error 5: I am proud that my Administration, with the support of the Legislature, is doing more to make significant improvements in mental health care than any since Governor Lurleen Wallace in the 1960s. Error reason: the model argues the “significant improvements” lack detail. Error 6: At times, her schoolwork and distance from her home state made her wonder if she should give up her Miss Alaska title. Error reason: the model argues the “significant improvements” lack detail. Error 7: It’s the result of great investment decisions, policies, vision, and direction. Error reason: not enough context about “it”. Error 8: Together with these partners, we’ll build a stronger, more durable health care system in Alaska that can respond to most any situation. Error reason: the model argues it lacks details about the plan. Error 9: At the same time, our ability to increase production is under attack from Washington, DC, and federal courts that side with extremist environmental groups. Error reason: the model argues it lacks details about the “attack”. Error 10: No state has been targeted more by the current administration than our Great State of Alaska. Error reason: the model argues it lacks details about specific actions. Error 11: No state has been targeted more by the current administration than our Great State of Alaska. Error reason: the model argues it lacks details about specific actions. Error 12: At every turn and since day one of the Biden Administration, this hostility has been perfectly clear. Error reason: the model argues it lacks details about the “hostility”. Error 13: Because no president should have to beg for more oil from the Middle East or Russia’s Arctic when we can produce it right here better and safer than anywhere else on the planet! This is common sense! Error reason: the model argues it lacks details or evidence about “the US produces better oil”. Error 13: Many of them have been with us for so long that they’ve almost been normalized in Alaska, as almost unsolvable. Error reason: lacking context and details about “long issues”. Error 14: I will always stand between Alaskans and a federal government that violates our God-given rights and exceeds its constitutional authority. Error reason: the model argues it lacks details about specific actions of the federal government and the speaker’s future action. Error 15: I envision an Alaska where our cost of energy is no longer the second-highest in the nation, but one of the lowest. That’s my vision. I hope it is yours as well. Error reason: the model argues it lacks details about the definition of “second-highest” and “lowest”. Error 16: I’ve seen it in the men and women on the frontlines of this pandemic who have helped us achieve one of the shortest shutdowns and one of the lowest death rates in the country. Error reason: the model argues it lacks details about “shortest” and “lowest”. Error 17: And because we want to lead by example, we are saving Coloradans money by making your State Government more efficient and effective. Error reason: the model argues it lacks details about “efficient” and “effective”. Error 18: Just as an earthquake is followed by aftershocks, we know that the overarching crisis of the pandemic has led to many other crises, perhaps lesser seen, but no less important to address. Error reason: the model argues it lacks details about “the crises”. Error 19: We owe it to the people of Colorado to improve safety and make Colorado truly one of the ten safest states in the nation over the next five years. Error reason: the model argues it lacks details about the speaker’s plan. Error 20: No other place offers opportunity to so many from such diverse backgrounds. Error reason: the model argues it lacks specific details. Error 21: It’s that, as our businesses grow, we don’t leave our workers behind. Error reason: the model argues it lacks specific details about business growth. Error 22: By creating choices - real choices - for parents, and unprecedented support for their kids. Error reason: the model argues it lacks specific details about the choices and supports. |
@@
| False Negative: understand facts as opinions or fail to identify facts entangled with opinions | Error 23: They plan their lives around hunting season, or fishing season; construction season, or tourism season.But not election season. Error reason: the model misunderstands it as the speaker’s opinion. But people’s lifestyles and priorities can be verified with related surveys or studies. Error 24: A future where a dynamic, multi-modal transportation system meets the needs of our growing population. Error reason: the model fails to identify “our growing population” as a fact. Error 25: When I was elected Governor, I knew that I would be remembered not for who I was, where I came from, or even what I said at events like this, but for what I did to make a meaningful, measurable, positive impact on the lives of Coloradans. Error reason: the model fails to identify that “the speaker is elected as the governor” is a fact. Error 26: But over time, we’ve learned we can’t solve big problems like climate change situationally, with short-term thinking. Error reason: the model fails to identify the causality claim about short-term thinking and big problems. Error 27: But at a time, when we’ve been heating and burning up, one thing we cannot do is repeat the mistakes of the past by embracing polluters. Error reason: the model fails to recognize the fact of embracing polluters in the past. |
| False Negative: hallucinate about personal experience and citation | Error 28: At times, her schoolwork and distance from her home state made her wonder if she should give up her Miss Alaska title. Error reason: the model argues the personal experience is subjective. Error 29: “A lot of people,” she said, “don’t recognize that their low points are what are going to propel them to their future. Error reason: subjective personal experience. Error 30: I agree with former Governor Jay Hammond that the government should never take more from the Permanent Fund than is distributed to the people of Alaska. Error reason: fail to detect the citation. Error 31: She is in a healthy marriage and is reconnecting with her children. Error reason: consider personal experience as unverifiable. Error 32: “Dad,” Catherine said, “Alaska has so much to offer.” Error reason: fail to detect the citation. Error 33: Still, she found the strength to take down the shooter, ending his violent killing spree and saving many precious lives. Error reason: consider personal experience as unverifiable. |
| False Negative: hallucinate about rhetoric | Error 34: They’re wondering how we’ve come to a place where the PFD is nothing more than what’s left over after government takes the lion’s share. Error reason: fail to understand the metaphor. |
@@
| False Positive: over-sensitive to granular, unspecific, or not-explicitly-presented facts | Error 1: Requesting to work from home because of the #coronavirus is what’s called a “reasonable accommodation.” You have disabled people to thank for that. Remember this moment in history the next time you think Accessibility laws are too “burdensome” to be abided. Error reason: the model recognizes the concept of “reasonable accommondation” and the existence of “accessibility laws” as facts, which are not explicitly presented by the post. |
| False Negative: misunderstand verifiable fact as subjective interpretation | Error 2: “Last week Trump told aides he’s afraid journalists will try to purposefully contract #coronavirus to give it to him on Air Force One.” https://t.co/sS1MZR6D7w Error reason: GPT-4 understands it as the tweet author’s subjective interpretation of Trump’s words. However, we think that it can be verified by checking whether Trump said the words or not. Error 3: Due to #coronavirus, media advises the economy must tank, the people must panic, Trump must be blamed, Biden must be secreted away from the public, and Bernie must cease rallies. I wonder why people do not trust the media’s motives on this? Error reason: GPT-4 understands it as the tweet author’s subjective interpretation of the media’s advice. However, we think it can be verified by checking if there are media suggesting such information. |
@@
| False Negative: fail to identify facts entangled with opinions | Error 1: Who would you prefer to lead our nation’s response to the growing #coronavirus threat? Error reason: fail to identify “the growing coronavirus threat”. Error 2: It was a really really really really really really really really really really really really really really really really really really really really bad idea to elect Donald Trump President of the United States. #TrumpVirus #TrumpCrash #TrumpRecession #COVID19 #coronavirus Error reason: fail to identify “elected Donald Trump Predisent of the US”. Error 3: If people who are infected by corona virus in SA were black, their names, homes street: pictures will be all over Social Media. White privileges goes a long way. Wait for a case of a black person, they will mention even his location they won’t say WC, they’ll say Gugulethu ext 5 Error reason: fail to identify the correlation between the infected persons’ race and their suffers. Error 4: @realDonaldTrump On a morning when Americans are terrified, the markets are gonna historically crash and we need LEADERSHIP…all you’ve done is hate-tweet BULLSHIT about Sanders, Warren, Biden, Democrats, Schumer. the media and now Obama. Your incompetence is staggering…. #Trump #coronavirus Error reason: although have subjective interpretations, the facts that “markets are gonna historically crash” and Trump commented something about others is verifiable. Error 5: Dear BBC, I want to fight for you. I know you’re more than news (which has been questionable) you’re also great drama, documentaries, kids tv etc But don’t make me question that by inviting Farage on to talk about Corona FFS!! Show you’ll fight for your integrity yourselves! Error reason: fail to identify the verifiable fact that BBC invited Farage is verifiable. Error 6: Thread 1: One wonders about the racial politics of this corona outbreak. What would have happened had it been blacks who came into the country with the virus? Would they hav been allowed to “self quarantine”? If the virus was from the continent; wouldn’t travels be banned by now? Error reason: fail to identify “this corona outbreak” and the practice of “self-quarantine”. Error 7: The total Iranian #COVID19 case-count is in the hundreds of thousands, perhaps millions, according to my estimates (detailed at the link). This raises an important question: if there are two million cases, where are all the bodies? https://t.co/nHYbQlXlVC Error reason: GPT-3.5 understands it as the tweet author’s subjective interpretation about the numbers. However, it can be verified by checking details in the link and reliable data source. Error 8: Due to #coronavirus, media advises the economy must tank, the people must panic, Trump must be blamed, Biden must be secreted away from the public, and Bernie must cease rallies. I wonder why people don’t trust the media’s motives on this? Error reason: GPT-3.5 understands it as the tweet author’s subjective interpretation about the media’s advice. However, we think it can be verified by checking if there are medias suggesting such information. |
@@
| False Negative: fail to comprehend claims about attached links | Error 9: Public Safety Announcement Fighting #CoronaVirus. We have to do this together. Wishing good health to all of you! Love, Vijay. https://t.co/fbafmmtq8S Error reason: this tweet claims that the link contains a public safetly announcement fighting COVID, which is verifiable. Error 10: This thread needs to fly. It shows how the legacy media is USING covid-19 as a political weapon and even how the SAME reporters are contradicting themselves. This. Is. SICK. https://t.co/Werq544xii Error reason: this tweet claims the content of the link, which is verifiable. |
@@
| False Negative: fail to identify personal experience or citation | Error 11: Beware the spread of coronavirus and Fox News pandemic propaganda. Trish Regan’s melodramatic rant decrying Dems and MSM for allegedly exploiting #COVID19 as “another attempt to impeach the President.” Yank this dangerous shrew off the air. #Trumpdemic https://t.co/6B60RLMIS0 Error reason: fail to identify the personal experience of Trish Regan. Error 12: I keep bumping into this problem. I want to be able to stand up and unequivocally defend the BBC. But it has repeatedly shut out radical voices and crucial issues while providing a massive platform for the alt-right to spout ill-informed nonsense. It is hard to love. https://t.co/A1pQMsqDxV Error reason: The interpretation about BBC’s behavior is subjective. But the tweet author’s previous stance might be verifiable by checking their previous statements. Error 13: @keywilliamss One African man actually got Corona and was cured in like a week. Health care officials are “baffled as to why Africa is virtually unscathed.” Which is… kinda racist that they expected it to be but lemme hush URL: https://t.co/yvP0DUXDiX Error reason: fail to identify the African man’s experience and the quotation of health care officials’ speech. |
@@
We conduct a thorough analysis on GPT-4 and GPT-3.5 AFaCTA. Errors on PoliClaimtest can be found in Table 8, Table 9, and Table 10. Errors on CheckThat!-2021-dev can be found in Table 11 and Table 12.